OpenStudy (anonymous):

What annual simple interest rate will allow Hosea Soli to increase his initial investment of $80,000 to $100,000 in five years?

OpenStudy (mathstudent55):

An increase from $80,000 to $100,000 is how much of an increase?

OpenStudy (anonymous):

20,000 @mathstudent55

OpenStudy (mathstudent55):

Great. This increase of $20,000 will take place over a 5 year period. How much of an increase is it per year?

OpenStudy (anonymous):

4000? @mathstudent55

OpenStudy (mathstudent55):

Correct again. One more step.

OpenStudy (mathstudent55):

Now you need to find: what percent of $80,000 is $4000?

OpenStudy (mathstudent55):

$80,000 is the whole so it corresoponds to 100%. $4000 is the part, so it corresponds to x, since we don't know it. We can use a proportion. 80000 is to 100% as 4000 is to x%. \( \dfrac{80000}{100} = \dfrac{4000}{x} \) Now you can cross multiply and solve for x.

OpenStudy (anonymous):

Okay. So I got 2. @mathstudent55

OpenStudy (mathstudent55):

2? 80000x = 4000 * 100 The answer is not 2.

OpenStudy (anonymous):

oops. I did 400 not 4000. @mathstudent55

OpenStudy (anonymous):

kay. Now i got 5 @mathstudent55

OpenStudy (mathstudent55):

That's better. That's correct. Now we can check. Find 5% of $80,000: 5% * $80,000 = 0.05 * $80,000 = $4,000 Multiply that amount by 5 (5 years): $4000 * 5 = $20,000 That is the increase that we calculated, so it's correct. The answer is 5%.

OpenStudy (anonymous):

Yay! You're awesome. Thank you!@mathstudent55

OpenStudy (mathstudent55):

Thanks. You're welcome.